### v1

@@ 5 Outlook and future developments
POSiCS demonstrated high spatial resolution when equipped with a LEHR collimator and showed good sensitivity in its LEHS collimator configuration. By comparing the results described in this work with current state-of-the-art IGCs, we demonstrate that POSiCS performances are comparable to, or superior to, currently available systems [Farnworth2023, Tsuchimochi_2013], while offering additional benefits, such as a lightweight module (less than 400 g) and wireless communication. These features significantly improve usability within the surgical environment. The most critical performance parameters determined through the tests described in this work are summarized in Table 4.
@@
Furthermore, the camera successfully detected gamma emissions from lutetium-177, demonstrating imaging capability up to the 113 keV photon emission line of this isotope. It was also possible to produce reliable images of the source by detecting lower-energy X-rays. These findings suggest that the device’s field of applications extends beyond traditional techniques involving the administration of Tc-99m, and may be suited for applications in theranostics and radiotherapeutic dosimetry, without the need for modifications to the detector system.
Acknowledgments
@@
This study was conducted as part of the POSiCS project. The POSiCS project is being carried out jointly by a consortium between University of Geneva (UNIGe), PIN lab at Hôpitaux Universitaires de Genève (HUG), and Fondazione Bruno Kessler (FBK) in Italy. POSiCS was funded for the second phase of the ATTRACT European initiative in June 2022, following the Horizon 2020 call. The Consortium Partners receive funding from the European Union for ATTRACT under Grant Agreement no. 101004462 and the Horizon 2020 Framework Programme for Research and Innovation (2014–2020). The position-sensitive LG-SiPM is based on the patented technology (EP3063559) developed by FBK.

### v2

@@ 5 Conclusions
POSiCS demonstrated high spatial resolution when equipped with a LEHR collimator and showed good sensitivity in its LEHS collimator configuration. By comparing the results described in this work with current state-of-the-art IGCs, we demonstrated that POSiCS performances are comparable to, or superior to, currently available systems [Farnworth2023, Tsuchimochi_2013], while offering additional benefits, such as a lightweight module (less than 400 g) and wireless communication. These features significantly improve usability within the surgical environment.
@@
Furthermore, the camera successfully detected gamma emissions from Lu-177, demonstrating imaging capability up to the 113 keV photon emission line of this isotope. It was also possible to produce reliable images of the source by detecting lower-energy X-rays. These findings suggest that the device’s field of applications extends beyond traditional techniques involving the administration of Tc-99m, and may be suited for applications in theranostics and radiotherapeutic dosimetry, without the need for modifications to the detector system.
List of abbreviations
| AEC | After energy cut |
| ADC | Analog-to-digital converter |
| BEC | Before energy cut |
| CZT | Cadmium–zinc–telluride |
| CoG | Center of gravity |
| DAC | Digital-to-analog converter |
| FOV | Field of view |
| FGS | Fluorescence-guided surgery |
| GATE | Geant4 Application for Tomographic Emission |
| HR | High-resolution (mode or collimator) |
| HS | High-sensitivity (mode or collimator) |
| IGC | Intraoperative gamma camera |
| LEHR | Low-energy high-resolution (collimator) |
| LEHS | Low-energy high-sensitivity (collimator) |
| LG-SiPM | Linearly graded silicon photomultiplier |
| LUT | Look-up table |
| LYSO:Ce | Lutetium–yttrium oxyorthosilicate, cerium-doped |
| NEMA | National Electrical Manufacturers Association |
| PET | Positron emission tomography |
| PMMA | Polymethyl methacrylate |
| PS | Position-sensitive |
| RGS | Radio-guided surgery |
| ROI | Region of interest |
| ROLL | Radioguided occult lesion localization |
| SC | Slow control |
| SPECT | Single-photon emission computed tomography |
| SiPM | Silicon photomultiplier |
| SLNB | Sentinel lymph node biopsy |
| SPAD | Single-photon avalanche diode |
| TIA | Transimpedance amplifier |
| UFOV | Useful field of view |
Declarations
Availability of data and material
The datasets used and/or analysed during the current study are available from the corresponding author on reasonable request.
Competing interests
The authors declare no competing interests.
Authors’ contributions
AR wrote the article, performed all tests, implemented the analysis code, and contributed to the image correction algorithm. FA provided the detector and wireless module, performed tests on the electronics, and revised the article. CA implemented the image reconstruction algorithm, contributed to the conceptualization, performed part of the tests, and revised the article. HA performed tests, contributed to the conceptualization, participated in the drafting, and revised the article. DdV supervised the entire project, contributed to the conceptualization, contributed to the drafting, and revised the article. AG provided resources, contributed to validation, and supported the hardware development. HZ supervised the work, contributed to the conceptualization, and revised the article. All authors read and approved the final manuscript.
Aknowledgements
@@
This study was conducted as part of the POSiCS project. POSiCS-2 was funded for the second phase of the ATTRACT European initiative in June 2022, following the Horizon 2020 call. The Consortium Partners receive funding from the European Union for ATTRACT under Grant Agreement no. 101004462 and the Horizon 2020 Framework Programme for Research and Innovation (2014–2020).
The POSiCS project is being carried out jointly by a consortium between University of Geneva (UNIGe) (Department of Nuclear and Corpuscolar Physics, PIN lab at Hôpitaux Universitaires de Genève (HUG), and Fondazione Bruno Kessler (FBK) in Italy. The position-sensitive LG-SiPM is based on the patented technology (EP3063559) developed by FBK.